GDP C I G AD IU 50 10 50 10 50 10 300 300 50 10...

GDP

C

I

G

AD

IU

50

10

50

10

50

10

300

300

50

10

400

380

50

10

500

460

50

10

600

540

50

10

700

620

50

10

680

20

Questions 1 – 4 are based on this table

1. Fill in the missing data in every empty cell above.                                                         20 pts

2.   Find the MPC and MPS for the above data. Show work!                                              10 pts

3.   Find the formula for Aggregate Demand using the above data and calculate equilibrium GDP.                                                                                                                                               20 pts

4.   Find the multiplier.   If full employment is 700 what is the necessary change in government purchases to move the economy to full employment? Show work!                                               20 pts
Answer #1

1) The table is completed using AD = C + I + G and IU = GDP - AD

GDP C I G AD IU
50 10 60
50 10 60
50 10 60
300 300 50 10 360 -60
400 380 50 10 440 -40
500 460 50 10 520 -20
600 540 50 10 600 0
700 620 50 10 680 20

b) MPC is the ratio of change in consumption to change in income. This is (380 - 300)/(400-300) = 0.80. MPS = 1 - MPC = 1 - 0.80 = 0.20

c) AD = C + I + G and here AD = Y = 600. This is the equilibrium income / GDP

d) The multiplier is 1/MPS = 1/0.20 = 5.   If full employment is 700 government needs to increase GDP by 100 (current GDP is 600). Hence the required increase in Government spending is (100/5) = 20